2 angles are supplementary. The measure of the smaller angle is 1/2 of the measure of the larger angle. Find the measure of the larger angle.
Answer by swincher4391(1107)

Let l +s =180
Also let s = 1/2l
2s = l
2s + s = 180
3s = 180
s = 60

Plug back in:
l + 60 = 180
l = 180 -60
l = 120.

60 and 120.